OpenStudy (anonymous):

(Please, show how you got your answer, I'm not sure how to do this...) Segment ST is congruent to segment SR and QR is congruent to segment SR. ___ ___ QR=x+4 and ST=5(3x-2) x+4=5(3x-2) -4 -4 x= 5(3x-6) Now.....what do I do now?

OpenStudy (anonymous):

you want to multiply out (or distribute) that 5 on the right hand side of the equation: \[5(3x-6) = 5*3x-5*6 = 15x-30\]

OpenStudy (anonymous):

5(3x-6)=5(3x)-5(6)=...something 15x-30=x !!!!! x=15x-30 -15x -15x?

OpenStudy (anonymous):

right. you are on the right track :)

OpenStudy (anonymous):

So... -14x=-30 ---- --- -14 -14? x=15/7?

OpenStudy (anonymous):

? Mahalo nui loa for the medal.... :)

OpenStudy (anonymous):

ack, i didnt notice what you had done in the very beginning was incorrect >.< by bad. Lets start from the beginning.

OpenStudy (anonymous):

Really? Ok. Let's do this!!!!

OpenStudy (anonymous):

you have: \[x+4 = 5(3x-2)\] From there, you need to multiply that 5 out first before you subtract 4 from both sides of the equation.

OpenStudy (anonymous):

x+4=5(3x-2)... So I divide 5 out of the equation first?

OpenStudy (anonymous):

you have to multiply it out, like this: \[5(3x-2) = 5*3x-5*2 = 15x-10\]

OpenStudy (anonymous):

15x-10=x+4 Now do I subtract 5?

OpenStudy (anonymous):

I mean 15x...

OpenStudy (anonymous):

you can do that if you want. That would leave us with: 15x - 10 = x + 4 -15x -15x -10 = -14x+4

OpenStudy (anonymous):

So now we are at: -10 = -14x + 4 what should we do next?

OpenStudy (anonymous):

Get rid of the four.... -14=-14x

OpenStudy (anonymous):

thats correct. now one more move....

OpenStudy (anonymous):

-14/-14=-14x/-14 1? x=1?

OpenStudy (anonymous):

yes, that answer looks a lot nicer than what we had before lol.

OpenStudy (anonymous):

:) Ok. Thank you for your help. I appreciate it! x)

OpenStudy (anonymous):

no problem :)